OpenStudy (anonymous):

evaluate f(x)= -sqrt2x+7 for f(37) a. -3 b. 0 c. -9 d. Not a real number

OpenStudy (anonymous):

\[-\sqrt{2x+7} for f(37)\]

OpenStudy (anonymous):

\[\Large \begin{array}{l} f(x) = - \sqrt {2x + 7} \\ f(37) = - \sqrt {(2 \times 37) + 7} \\ = - \sqrt {81} \\ = - 9 \end{array}\]